How much power is exerted to lift a rock weighing 11.9 kg to a height of 1.23 m in 1.31 s?


1
Expert's answer
2021-04-28T09:27:02-0400

Total force exerted by rock "=mg=11.9\times9.8=116.62\space N"

Total Displacement of rock, "s=1.23\space m"

Work Done "=Force\times Displacement=116.62\times1.23=143.44\space J"

Time, "t=1.31\space s"

Power "=\dfrac{Work \space Done}{Time}=\dfrac{143.44}{1.31}=109.49\space W"